# Horizon-Length Prediction: Advancing Fill-in-the-Middle Capabilities for Code Generation with Lookahead Planning

### v2

@@ 4.2 Repository-Level Cross-File Code Fill-in-the-Middle
4.3 Code Fixing via Fill-in-the-Middle
We use Defects4J (Just et al., 2014) to evaluate the performance of HLP on code fixing. Defects4J consists of open-source bugs found across 15 Java repositories. Following existing works (Xia et al., 2023; Xia & Zhang, 2023), we collect 313 single-hunk bugs from Defects4J that can be fixed by replacing or adding a continuous code hunk. Specifically, for each bug, models are prompted to generate the correct code hunk (i.e., patch) given the left and right contexts of the buggy code hunk, and the whole test suite of the project will be executed to evaluate the correctness of the generated patch. Patches that can successfully pass the test suite are referred to as plausible patches and we report the number of plausible patches as our evaluation metric. As shown in the "Code Fix" section of Table 5, adding HLP during training results in relatively up to 18% more bugs fixed by the model222Note that Defects4J is a small dataset with only 313 examples and models can only solve 30-60 out of those, which makes it hard to obtain statistically significant differences with greedy decoding..
| Code Fix | Code Reasoning | ||
@@ 4.4 Code Reasoning via Fill-in-the-Middle
Lastly, we examine whether HLP improves model performance on code reasoning beyond ordinary completion use cases. The motivation behind is that HLP aims to teach the model to plan ahead, and planning is a special subset of reasoning that requires an action sequence over a long time horizon (Kang et al., 2024). Towards this end, we consider CRUXEval (Gu et al., 2024) which comprises 800 Python functions paired with two distinct tasks: CRUXEval-I, where LLMs need to predict the input from the known output, and CRUXEval-O, where LLMs are required to predict the output based on the given input.
We reformat prompts of CRUXEval-I into FIM style and leave CRUXEval-O as L2R generation, both of which are evaluated in zero-shot setting. Different from previous subsections where post-processing is not used, we follow the same pipeline as in the original CRUXEval paper to extract accurate input/output values from generation because we are focusing on evaluating the reasoning capability of LLMs rather than their capability of generating correct code333In CRUXEval-I, we only want to evaluate the correctness of the input value infilled by LLMs in the given assertion. However, FIM-style prompts we use in the experiments does not restrict LLMs from writing multiple assertions before starting infilling the given assertion, which is useless in this task. So we use post-processing techniques to extract the input value infilled for the given assertion to better evaluate the reasoning capabilities.. As shown in the "Code Reasoning" section of Table 5, HLP demonstrates up to 6% improvements on both CRUXEval-I and CRUXEval-O tasks for all the code LLMs consistently, which shows that HLP also improves code reasoning capabilities of LLMs.
@@ 5 Discussion
5.1 Comparing HLP with Multi-token Prediction
@@
| Algorithmic | Algorithmic | Control | API | ||
@@
5.2 Next-Token Prediction Alone Cannot Yield Horizon Awareness
| Training | Test | |
| DS-1.3B | 0.455 | 0.440 |
@@ 5.2 Next-Token Prediction Alone Cannot Yield Horizon Awareness
| DS-6.7B | 0.525 | 0.519 |
[…]
| SC2-3B | 0.364 | 0.356 |
[…]
| SC2-7B | 0.418 | 0.410 |
[…]
We argue that NTP alone does not grant code LLMs the awareness of prediction horizon. Specifically, we show that hidden states of baseline models trained with NTP only do not carry information about the number of future tokens (i.e., horizon length). Consequently, including HLP in training is essential for models to be knowledgeable of prediction horizon.
We design a probing task by fitting linear regression models over hidden states of code LLMs trained with or without HLP respectively, while freezing all parameters of the underlying transformers. In our experiments, we use different models to generate hidden states for 20K code snippets, which gives hidden state vectors for 7.8M tokens from the part. We split them into the training and the test set, ensuring no overlap at sequence level. Taking these hidden state vectors as inputs and the true normalized remaining token counts as targets, we fit two linear regression models for each code LLM trained with or without HLP, respectively.
We plot the predicted percentage of remaining tokens versus the normalized token position in Figure 4, and report the coefficient of determination in Table 7. As shown, the regression model does not fit well with hidden states from the baseline model, indicating that those hidden states are not strongly correlated with future token count and do not contain information about horizon length. In contrast, with HLP, the hidden states perform much better. The result demonstrates that horizon awareness does not naturally exist in language models trained with NTP, and can only be obtained through targeted training tasks such as HLP.
6 Related Work
Fill-in-the-Middle for Code Language Models The unprecedented success of causal language models such as GPT-3 (Brown et al., 2020) in natural language has inspired researchers to develop similar decoder-only models for programming languages. These models are trained on massive source code data for applications such as code generation. While early models such as Codex (Chen et al., 2021) and CodeGen (Nijkamp et al., 2023) only support Left-to-Right (L2R) generation, Fill-in-the-Middle (or infilling) has attracted increased attention because right context naturally carries an indispensable part of information for completing code in the middle (Fried et al., 2023; Bavarian et al., 2022). Subsequently, FIM training has become a common practice widely adopted by most code LLMs, such as StarCoder (Li et al., 2023; Lozhkov et al., 2024), DeepSeek-Coder (Guo et al., 2024; DeepSeek-AI et al., 2024), and Code Llama (Rozière et al., 2023).
Existing models generally tackle the infilling problem by breaking a code snippet into --, and reordering them into -- (PSM) or -- (SPM). The transformed sequences are fed to the model during training for standard next-token prediction (NTP). We point out that the infilling task cannot be effectively learned with NTP alone, as it requires planning capability for the model to fluently and meaningfully connect the middle completion to the suffix through forward looking during auto-regressive decoding.
Apart from those, Gloeckle et al. (2024) proposed to predict multiple tokens from a single hidden state during both training and inference, which was shown to achieve stronger performance on coding tasks with no computation overhead. While multi-token prediction enhances models’ planning capability within the tokens predicted together (), we argue that with a small , the limited horizon is usually insufficient for planning in the case of infilling as the connection from to only happens towards the end of the generation. In contrast, HLP adopts a global and arbitrary long horizon over all future tokens by counting the remaining generation budget, which more effectively helps models to close the generation fluently with early planning.
7 Conclusion
Fill-in-the-Middle is ubiquitous in code completion, and therefore, has become an important consideration in the development of code language models. The current FIM training paradigm splits and reorders original training sequences (--) into FIM-style sequences (--/PSM or --/SPM), and performs standard next-token prediction. However, this approach frequently results in models struggling to generate content that smoothly aligns with the right context. While existing FIM benchmarks frequently rely on different post-processing methods to circumvent this problem, we emphasize that such methods typically require dataset-specific assumptions, which are impractical in real-world scenarios. To address this limitation and enhance the infilling capability of code language models, we propose Horizon-Length Prediction (HLP). HLP teaches models to predict the portion of remaining tokens at every step. Experiments across different model families and sizes show that HLP improves infilling performance on diverse FIM benchmarks, across file-level and repository-level, and without using any dataset-specific post-processing. Moreover, the enhanced planning capability acquired through HLP training also boosts models’ performance on code reasoning tasks, suggesting that HLP may broadly improve language models’ reasoning capabilities. Besides, HLP is also efficient as it does not cause any inference overhead and the training overhead is negligible as well. Our work marks a significant advancement in developing more effective code language models for real-world applications.
Appendix A Appendix
A.1 Effect of HLP on Left-to-Right Performance
@@
Therefore, instead of taking the entire code file as prediction horizon, we ask the model to predict the number of future tokens required to complete current line in L2R training, which is a natural semantic unit in code. Furthermore, to avoid conflicts between and , we use two independent s to let the model learn and separately. As shown in Table 8, by applying and simultaneously, the performance degradation on L2R tasks is recovered, with the improvement on FIM tasks largely retained. These results demonstrate the generalizable effectiveness of HLP and shows the huge potential of applying the idea of HLP to more general training scenarios.
A.2 Ablation Studies
@@
| Algorithmic | Algorithmic | Control | API | ||
@@
Applying HLP to all tokens v.s. first token only. While it is easy to see that knowing the HLP loss on the first token is sufficient to infer the horizon length in theory, having HLP loss on every token provides denser and consistent supervision signals which makes learning easier (as discussed in §3). It also helps regularize the hidden representation of every subsequent token. To empirically show this, we conduct an experiment by only applying HLP to the first token only (i.e., “HLP (first)”) and compared its performance with our original HLP design (i.e., “HLP (all)”). As shown in Table 10, while applying HLP only to the first token performs better than NTP only, applying HLP loss for each token can achieve better performance than applying it to just the first token.
@@
| Algorithmic | Algorithmic | Control | API | ||
@@
Normalized v.s. unnormalized targets. We use normalized targets in our original HLP design (i.e., using rather than ) is that the scale of HLP loss will be otherwise in a huge range, e.g., some examples has single digit loss while some might have thousands. To further study the effect of normalization, we conduct an experiment by using as the target (i.e., "HLP (raw)") rather than (i.e., "HLP (normalized)"). To achieve this, we remove the sigmoid function from the original HLP. As shown in Table 11, setting the target as fails to improve FIM performance, likely due to the large-scale HLP loss after using as the target interferes with NTP pre-training.
@@
| Algorithmic | Algorithmic | Control | API | ||

### v3

@@ 4.2 Repository-Level Cross-File Code FIM
| Code Repair | Code Reasoning | ||
@@
4.3 Code Repair via Fill-in-the-Middle
To assess HLP’s impact on practical applications beyond code completion, we evaluate the performance of code repair using Defects4J (Just et al., 2014). Defects4J consists of open-source bugs found across 15 Java repositories. Following existing works (Xia et al., 2023; Xia and Zhang, 2023), we collect 313 single-hunk bugs from Defects4J that can be fixed by replacing or adding a continuous code hunk. Specifically, for each bug, models are prompted to generate the correct code hunk (i.e., patch) given the left and right contexts of the buggy code hunk, with correctness verified by project test suites. As shown in the “Code Repair” section of Table 5, adding HLP during training results in relatively up to 18% more bugs fixed by the model222Note that Defects4J is a small dataset with only 313 examples and models can only solve 30-60 out of those, which makes it hard to obtain statistically significant differences with greedy decoding., suggesting that enhanced planning capabilities translate directly to better bug-fixing performance.
@@ 4.4 Code Reasoning via Fill-in-the-Middle
We further examine whether HLP’s planning benefits extend to code reasoning tasks. We use CRUXEval (Gu et al., 2024) which comprises 800 Python functions paired with two distinct tasks: CRUXEval-I, where LLMs need to predict the input from the known output, and CRUXEval-O, where LLMs are required to predict the output based on the given input.
We reformat prompts of CRUXEval-I into FIM style and leave CRUXEval-O as L2R generation, both of which are evaluated in zero-shot setting. Different from previous subsections where post-processing is not used, we follow the same pipeline as in the original CRUXEval paper to extract accurate input/output values from generation because we are focusing on evaluating the reasoning capability of LLMs rather than their capability of generating correct code333In CRUXEval-I, we only want to evaluate the correctness of the input value infilled by LLMs in the given assertion. However, FIM-style prompts we use in the experiments does not restrict LLMs from writing multiple assertions before starting infilling the given assertion, which is useless in this task. So we use post-processing techniques to extract the input value infilled for the given assertion to better evaluate the reasoning capabilities.. As shown in the “Code Reasoning” section of Table 5, HLP demonstrates up to 6% improvements on both CRUXEval-I and CRUXEval-O tasks for all the code LLMs consistently, which shows that HLP also improves intrinsic code reasoning capabilities of LLMs.
@@ 5 Discussion
5.1 NTP Alone Cannot Yield Horizon Awareness
A key question of interest is whether standard next-token prediction (NTP) inherently provides models with awareness of generation horizons. Through systematic analysis, we demonstrate that hidden states of models trained with NTP alone do not capture meaningful information about required completion lengths.
To quantify this, we design a probing task that attempts to predict the remaining completion length from s’ hidden states. We extract hidden states from 20K code snippets (approximately 7.8M tokens) using models trained with and without HLP and split the data to ensure no sequence-level overlap between train and test. Using these hidden states as input and normalized remaining token counts as targets, we fit linear regression models while keeping the underlying transformer parameters frozen.
| Test | DS-1.3B | DS-6.7B | SC2-3B | SC2-7B |
| NTP | 0.440 | 0.519 | 0.356 | 0.410 |
[…]
Figure 3 shows predicted versus actual remaining token percentages at different positions, while Table 6 reports the coefficient of determination . Models trained with NTP alone show poor fit, indicating their hidden states lack horizon information. In contrast, models trained with HLP show much stronger correlations between hidden states and remaining lengths. This result demonstrates that horizon awareness must be explicitly trained rather than emerging naturally from NTP. We also carried out a non-linear probing experiment with the same inputs and targets by replacing linear regression models with an MLP regressor that uses logistic activation function (Appendix A.2), which further confirms our finding here.
5.2 Why Horizon-Length Prediction works?
The effectiveness of Horizon-Length Prediction (HLP) stems from its ability to address a critical limitation of standard next-token prediction (NTP) in autoregressive models: the absence of global planning awareness. While NTP optimizes for local token-level coherence by maximizing the likelihood of immediate next tokens, it inherently struggles to enforce long-horizon structural consistency between the generated and the given in FIM. This limitation becomes pronounced when the model must align generated code with both preceding logic and subsequent constraints (e.g., API calls or variable dependencies defined in the suffix).
HLP bridges this gap by teaching models to explicitly condition the model on the remaining generation horizon at every decoding step. By regressing the normalized count of future tokens required to complete , HLP effectively decomposes the infilling task into a sequence of length-aware subgoals, where each generated token is tied to a dynamically updated “budget” of remaining steps. This mechanism mirrors human problem-solving strategies, where progress estimation (e.g., "30% of steps remaining") guides iterative refinement of intermediate actions to meet global objectives.
The success of HLP is rooted in its complementary role to NTP. While NTP ensures local fluency, HLP provides a global scaffold for planning. By unifying local token prediction with global horizon awareness, HLP bridges the gap between autoregressive decoding and holistic reasoning, enabling models to dynamically adapt their generation strategy to long-horizon constraints. This synergy is particularly critical in code infilling, where the correctness of depends on both preceding logic (i.e., ) and subsequent context (i.e., ), demanding a balance between immediate token likelihood followed by and forward-looking structural coherence constrained by distant .
Our in-depth attention analysis provides direct evidence of this enhanced planning capability. Figure 4 shows that models trained with HLP pay significantly more attention to context, particularly at the start of the generation. This indicates that HLP enables proactive planning by considering long-horizon constraints from the beginning. Furthermore, the consistent improvements across FIM, bug-fixing, and reasoning tasks demonstrate that this enhanced planning capability generalizes beyond simple completion scenarios.
In summary, HLP transforms code generation from a myopic token-by-token process into a structured planning task. It enables models to dynamically adapt their generation strategy to long-horizon constraints, ensuring that each step not only satisfies local fluency but also contributes to a globally coherent solution. This dual focus on immediate actions and overarching goals positions HLP as a pivotal advancement in code LLMs.
5.3 HLP Mitigates Planning Failures in FIM
We conduct a quantitative analysis to empirically demonstrate the frequency of planning failures in FIM and evaluate the effectiveness of HLP in addressing these failures. As presented in Figure 1 (§2), a prevalent manifestation of planning failures in FIM is the "correct code but incorrect order" pattern, wherein the model generates the final line of ground truth code accurately but positions it incorrectly within the completion. To maintain analytical tractability, our study focuses specifically on planning failures that result in this particular pattern. We establish the following criteria to identify instances of the "correct code but incorrect order" pattern in model’s FIM completions:
Completion fails to pass all the unit tests.
Completion contains the last line from ground truth.
Completion does not end with this last line.
Testing on SAFIM’s Algorithmic and Algorithmic_v2 partitions which focus on multi-line FIM, we find that models trained without HLP exhibit the “correct code but incorrect order” pattern in 5-10% of problems across the benchmark, as shown in Table 7. With HLP training, this drops to 4-6%, demonstrating that HLP helps alleviate lack-of-planning failures effectively.
| SAFIM | Avg | ||
| Algo | Algo | ||
| DS-1.3B | 9.34% | 10.48% | 9.91% |
[…]
| DS-6.7B | 5.16% | 5.93% | 5.55% |
[…]
| SC2-3B | 7.25% | 7.35% | 7.30% |
[…]
| SC2-7B | 5.79% | 6.26% | 6.03% |
[…]
6 Related Work
Fill-in-the-Middle for Code Language Models Large language models trained on massive source code data have demonstrated great potential in various applications for software development. While early models such as Codex (Chen et al., 2021) and CodeGen (Nijkamp et al., 2023) only support Left-to-Right (L2R) generation, Fill-in-the-Middle (or infilling) has attracted increased attention because right context naturally carries an indispensable part of information for completing code in the middle (Fried et al., 2023; Bavarian et al., 2022). Subsequently, FIM training has become a common practice widely adopted by most code LLMs, such as StarCoder (Li et al., 2023; Lozhkov et al., 2024), DeepSeek-Coder (Guo et al., 2024; DeepSeek-AI et al., 2024), and Code Llama (Rozière et al., 2023).
Existing models generally tackle the infilling problem by breaking a code snippet into --, and reordering them into -- (PSM) or -- (SPM), which are then used for next-token prediction (NTP) training. We point out that the infilling task cannot be effectively learned with NTP alone, as it requires planning capability for the model to fluently and meaningfully connect the middle completion to through forward looking during auto-regressive decoding.
Apart from those, Gloeckle et al. (2024) proposed to predict multiple tokens from a single hidden state during both training and inference, which was shown to achieve stronger performance on coding tasks with no computation overhead. While multi-token prediction enhances models’ planning capability within the tokens predicted together (), we argue that with a small , the limited horizon is usually insufficient for planning in the case of infilling as the connection from to only happens towards the end of the generation. In contrast, HLP adopts a global and arbitrary long horizon over all future tokens by counting the remaining generation budget, which more effectively helps models to close the generation fluently with early planning444Please refer to Appendix §A.3 for more details..
7 Conclusion and Future Work
Fill-in-the-Middle is ubiquitous in code completion, reflecting the iterative nature of software development where code is frequently inserted between existing sections, and thus has become an important consideration in the development of code language models. The current FIM training paradigm splits and reorders training sequences for next-token prediction. However, this approach frequently results in models struggling to generate content that smoothly aligns with the right context. While existing FIM benchmarks rely on different post-processing methods to circumvent this problem, we emphasize that such methods typically require dataset-specific assumptions, which are impractical in real-world scenarios.
To address this limitation and enhance the infilling capability of code language models, we propose Horizon-Length Prediction (HLP). HLP teaches models to predict the portion of remaining tokens at every step, enabling them to develop planning awareness over arbitrarily long horizons. Experiments across different model families and sizes show that HLP improves infilling performance on diverse FIM benchmarks, across file-level and repository-level, and without using any dataset-specific post-processing. Moreover, the enhanced planning capability acquired through HLP training also boosts models’ performance on code reasoning tasks, suggesting broader benefits for language models’ reasoning capabilities. Importantly, HLP achieves these improvements while maintaining efficiency, with negligible training overhead and no additional inference cost.
Our work marks a significant advancement in developing more effective code language models for real-world applications. Future work could explore extending HLP to other generation tasks requiring long-horizon planning and investigate its potential for enhancing general reasoning capabilities.
Limitations
While HLP has proven effective through extensive evaluation in the paper, our pre-training experiments are restricted to models of 7B parameters due to the computation budget. It is prohibitively expensive to perform pretraining experiments for LLM, and unfortunately we do not have enough resources to demonstrate the impact of HLP on larger models. This highlights a broader challenge in open science for pre-training research. In addition, this work has mainly focused on the code domain, but the idea of horizon-length prediction can be broadly applicable for improving models’ reasoning capability in natural language and mathematical domains as well, which we leave to future work.
Appendix A Appendix
A.1 Effect of Value Selection
Following the original experimental settings (but with 100k training steps given limited computational resources), we studied a series of values and report both HLP validation loss and SAFIM pass@1. As shown in Table 8, our method is robust to hyperparameter selection: all non-zero values improve over the baseline (). Moreover, we observe that HLP loss quickly decreases within the first few thousand training steps with all positive ’s, indicating that a large isn’t necessary for models to learn HLP well. This also coincides with our observation on downstream performance: while lower HLP loss generally correlates with better SAFIM performance, the improvement plateaus beyond a certain threshold, suggesting that once the model acquires basic horizon awareness, further optimization of HLP loss provides diminishing returns for downstream performance.
| HLP Valid Loss | SAFIM (pass@1) | |
| 0 | 0.254 | 47.40% |
| 0.01 | 0.060 | 48.98% |
| 0.1 | 0.048 | 49.10% |
| 0.2 | 0.045 | 48.74% |
| 0.5 | 0.043 | 49.34% |
A.2 Non-Linear Probing for Horizon Awareness
By replacing the linear regression models in the original linear probing experiment with an MLP regressor that uses logistic activation function, we further conduct a non-linear probing experiment to study whether NTP is able to yield horizon awareness. As shown in Table 9, even under the assumption of the non-linear relationship, models trained without HLP cannot effectively extract horizon information from their hidden states. This demonstrates that horizon awareness must be explicitly trained rather than emerging naturally from NTP.
| Test | DS-1.3B | DS-6.7B | SC2-3B | SC2-7B |
| NTP | 0.392 | 0.431 | 0.180 | 0.226 |
[…]
A.3 Comparing HLP with Multi-token Prediction
@@
| Algo | Algo | Control | API | ||
@@
A.4 HLP Improves Self-Infilling Performance
Recently, some strategies have been proposed to conduct L2R code generation based on FIM capabilities and self-infilling (Zheng et al., 2024) is one of the representative strategies of this kind. Given an initial input prompt , self-infilling performs L2R code generation by using FIM decoding to (1) generate based on and then (2) generate based on and previously generated . Furthermore, self-infilling proposes a looping mechanism to improve the generated code iteratively, where it first uses L2R decoding to generate a new ′ based on and previously generated and then uses FIM decoding to generate a new ′ based on and this newly generated ′. This looping procedure can be continued for multiple rounds to obtain greater improvements.
| HumanEval | |||||
| DS-1.3B | 27.4 | 29.3 | 31.7 | 31.7 | 31.7 |
@@ A.4 HLP Improves Self-Infilling Performance
Since self-infilling relies on FIM capabilities for L2R code generation, it is interesting to study whether HLP can also enhance its performance. To this end, we evaluate the self-infilling performance of DeepSeek-Coder-Base 1.3B trained with and without HLP on HumanEval using greedy decoding, following the setting of the original paper (Zheng et al., 2024). We report models’ performance on HumanEval with ranging from 0 to 4, where denotes the number of times the decoding process goes through the loop and represents that the looping mechanism is not activated. As shown in Table 11, the model trained with HLP consistently outperforms the model trained without HLP across different number of self-improving steps. Furthermore, with more self-improving steps, while the performance of the model trained without HLP gets stuck, the performance of the model trained with HLP continues to showcase steady improvements.
A.5 Effect of HLP on Left-to-Right Performance
@@
| Algo | Algo | Control | API | ||
@@
| Algo | Algo | Control | API | ||
@@
| Algo | Algo | Control | API | ||
@@
Therefore, instead of taking the entire code file as the prediction horizon, we ask the model to predict the number of future tokens required to complete current line in L2R training, which is a natural semantic unit in code. Furthermore, to avoid conflicts between and , we use two independent s to let the model learn and separately. As shown in Table 12, by applying and simultaneously, the performance degradation on L2R tasks is recovered, with the improvement on FIM tasks largely retained. These results demonstrate the generalizable effectiveness of HLP and shows the huge potential of applying the idea of HLP to more general training scenarios.
A.6 Additional Ablation Studies
Applying HLP to all tokens v.s. first token only. While it is easy to see that knowing the HLP loss on the first token is sufficient to infer the horizon length in theory, having HLP loss on every token provides denser and more consistent supervision signals which makes learning easier (as discussed in §3). It also helps regularize the hidden representation of every subsequent token. To empirically show this, we conduct an experiment by applying HLP to the first token only (i.e., “HLP (first)”) and compared its performance with our original HLP design (i.e., “HLP (all)”). As shown in Table 14, while applying HLP only to the first token performs better than NTP only, applying HLP loss for each token can achieve better performance than applying it to just the first token.
Normalized v.s. unnormalized targets. We use normalized targets in our original HLP design (i.e., using rather than ) is that the scale of HLP loss will be otherwise in a huge range, e.g., some examples has single digit loss while some might have thousands. To further study the effect of normalization, we conduct an experiment by using as the target (i.e., “HLP (raw)”) rather than (i.e., “HLP (normalized)”). To achieve this, we remove the sigmoid function from the original HLP. As shown in Table 15, setting the target as fails to improve FIM performance, likely due to the large-scale HLP loss after using as the target interferes with NTP pre-training.